OpenStudy (anonymous):

A wooord problem... I need some help Bocce is a lawn bowling game that originated in Italy. The bocce court is a rectangle with the artea of 1032square feet. The width of the court is 12 feet. what is the length of the court?

OpenStudy (anonymous):

area of rectangle = length * width 1032 = length * 12 1032/12 = length so length= 1032/12 = 86 feet ans